OpenStudy (anonymous):

The vertices of triangle IJK have coordinates I(−5, −2), J(−5, −4), and K(−4, −4). Which coordinates are the coordinates of the vertices of triangle IꞌJꞌKꞌ when triangle IJK is rotated 180° clockwise around the origin?

rishavraj (rishavraj):

see i if O is origin .....let P (x , y) with respect to axes OX and OY.....and let P(x' , y') with respect to axes OX' and OY' where angle X'OX = angle Y'OY = \[\theta \] then x' = x cos(theta) + y sin(theta) y' = -x sin(theta) + y cos(theta)

OpenStudy (anonymous):

okay then what ? @rishavraj

rishavraj (rishavraj):

u got the angle theta = 180 degrees.... so just plug the values

rishavraj (rishavraj):

like for I I' (x' , y') x' = -5 cos(180) + (-2) sin(180) y' = 5 sin(180) + (-2) cos(180) @princess_hannahleigh

OpenStudy (anonymous):

I still dont get it !

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

@princess_hannahleigh

rishavraj (rishavraj):

@princess_hannahleigh u there???